# Nyström Kernel Stein Discrepancy

### v1

@@ A.3 Proof of Corollary 1
    where (a) uses the exponential decay assumption. (b) uses that $n\geq m$ and that the logarithm is a monotonically increasing function. (c) follows from our choice of
    $m\gtrsim\sqrt{n}\sqrt{\log\left(1+\frac{c_{1}n}{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\right)\log(12n/\delta)\log(12/\delta)},$ (A.32)
    finishing the proof of the corollary.
@@ Appendix B Auxiliary results
This section collects our auxiliary results. Lemma B.1 builds on Rudi et al. [2015, Lemma 6], which assumes bounded feature maps, and on Della Vecchia et al. [2021, Lemma 5], which is stated in the context of leverage scores. Lemma B.2 states that a sub-exponential random variable satisfies Bernstein’s conditions, and Lemma B.3 shows that a “whitened” sub-Gaussian vector remains sub-Gaussian. Lemma B.4 expresses the $L_{2}$-norm of the RKHS-norm as trace. In Lemma B.5, we show how tensor products interplay with linearly transformed vectors. Lemma B.6 is about the maximum of real-valued sub-Gaussian random variables; it is a slightly altered restatement of Canonne [2021].
@@
Let Assumption 1 hold, and assume $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$. Then, for any $\delta\in(0,1)$, it holds with probability at least $1-\delta$ that
$\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\lesssim\lambda$ (B.33)
@@
The next lemma shows that the “whitened” random variable $C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)$ inherits Assumption 1(ii) from $h_{p}\left(\cdot,{X}\right)$.
Lemma B.3.
Let Assumption 1(ii) hold. Let $X\sim\mathbb{Q}$, $C_{\mathbb{Q},h_{p}}$ be the covariance operator, $C_{\mathbb{Q},h_{p},\lambda}=C_{\mathbb{Q},h_{p}}+\lambda I$, and $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$. Then, there exists an absolute constant $c>0$ such that
$\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(\dagger)}}{{\leq}}c\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}\stackrel{{\scriptstyle(\ddagger)}}{{<}}\infty.$ (B.50)
@@
We will refer to the $\leq$’ in the statement by ()($\dagger$)( † ), and to the ’$<$’ by ()($\ddagger$)( ‡ ). First, we observe that by the imposed Assumption 1(ii), there exists a constant $c_{1}>0$ such that
$\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\leq c_{1}\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty.$ (B.51)
which, by the definition of the $\left\|{\cdot}\right\|_{\psi_{2}}$-norm, implies that
$\mathbb{E}_{X\sim\mathbb{Q}}\exp\left(\frac{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}}{c_{1}^{2}\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}}\right)\leq 2.$ (B.52)
Our goal for ()($\dagger$)( † ) is to prove that there exists $c>0$ such that
$\mathbb{E}_{X\sim\mathbb{Q}}\exp\left(f(X;c)\right)$ $\leq 2,\quad\text{with}$ $f(X;c)$ $:=\frac{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}}{c^{2}\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}},$
and we will show ()($\ddagger$)( ‡ ) along the way. We bound the terms in the numerator and the denominator of $f(X;c)$ separately, and choose $c>0$ at the end.
  • Upper bound on $\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}$, and $\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$:
@@
    $\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}^{2}\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}\stackrel{{\scriptstyle(b)}}{{=}}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1}}\right\|_{\mathrm{op}}\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}.$ (B.53)
    (a) is implied by $\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}\leq\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ following from the definition of the operator norm, (b) comes from $\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}^{2}=\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1}}\right\|_{\mathrm{op}}$ holding by the $C^{*}$ property and the self-adjointness of $C_{\mathbb{Q},h_{p},\lambda}^{-1/2}$. Specifically, we got that
    $\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\left\|{\sqrt{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1}}\right\|_{\mathrm{op}}}\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}$
    =$\stackrel{{\scriptstyle(b)}}{{=}}\sqrt{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1}}\right\|_{\mathrm{op}}}\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}\stackrel{{\scriptstyle(c)}}{{<}}{\infty},$
    where (a) holds by (B.53) and the monotonicity of the integration, (b) is implied by the homogeneity of norms. (c), i.e., ()($\ddagger$)( ‡ ), holds since $\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$ by our assumption recalled in (B.51).
  • Lower bound on $\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}$:
    $\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)\stackrel{{\scriptstyle(b)}}{{=}}\sum_{i\in I}\frac{\lambda_{i}}{\lambda_{i}+\lambda}\stackrel{{\scriptstyle(c)}}{{\geq}}\sum_{i\in I}\frac{\lambda_{i}}{2\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}}$ (B.54)
    =$\stackrel{{\scriptstyle(d)}}{{=}}\frac{\operatorname{tr}(C_{\mathbb{Q},h_{p}})}{2\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}}.$ (B.55)
    (a) is stated in Lemma B.4. (b) holds by the eigenvalue-based form of the $\operatorname{tr}$, where $\left(\lambda_{i}\right)_{i\in I}$ are the eigenvalues of $C_{\mathbb{Q},h_{p}}$. $(c)$ is implied by $\lambda_{i}\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$ holding for all $\left(\lambda_{i}\right)_{i\in I}$ by the compact positive property of $C_{\mathbb{Q},h_{p}}$ and the assumption $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$. $(d)$ follows again from the eigenvalue-based form of the $\operatorname{tr}$.
    One can now rewrite $\operatorname{tr}(C_{\mathbb{Q},h_{p}})$ in the resulting expression (B.55) as
    $\operatorname{tr}(C_{\mathbb{Q},h_{p}})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\operatorname{tr}\left[\int_{\mathbb{R}^{d}}h_{p}\left(\cdot,{\mathbf{x}}\right)\otimes h_{p}\left(\cdot,{\mathbf{x}}\right)\mathrm{d}\mathbb{Q}(\mathbf{x})\right]\stackrel{{\scriptstyle(b)}}{{=}}\int_{\mathbb{R}^{d}}\operatorname{tr}\left[h_{p}\left(\cdot,{\mathbf{x}}\right)\otimes h_{p}\left(\cdot,{\mathbf{x}}\right)\right]\mathrm{d}\mathbb{Q}(\mathbf{x})$
@@
    =$\stackrel{{\scriptstyle(c)}}{{=}}\int_{\mathbb{R}^{d}}\left\|{h_{p}\left(\cdot,{\mathbf{x}}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}\mathrm{d}\mathbb{Q}(\mathbf{x})\stackrel{{\scriptstyle(d)}}{{=}}\left\|{\left\|{h_{p}\left(\cdot,{X}\right))}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}.$
    In (a) we used the definition of $C_{\mathbb{Q},h_{p}}$, in (b) we flipped the integration and the $\operatorname{tr}$ [Steinwart and Christmann, 2008, (A.32)]. (c) follows from Lemma D.1 by choosing $f:=h_{p}\left(\cdot,{\mathbf{x}}\right)$, in (d) we used the definition of $\left\|{\cdot}\right\|_{L_{2}(\mathbb{Q})}$.
    Combining this form with (B.55), we got the lower bound
    $\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}$ $\geq\frac{\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}}{2\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}}.$ (B.56)
Taking into account the upper bound (B.53) and the lower bound (B.56), we arrived at
$f(X;c)$ $\leq\frac{2\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1}}\right\|_{\mathrm{op}}\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}}{c^{2}\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{k}^{d}}}\right\|_{L_{2}(\mathbb{Q})}^{2}},$
which by the monotonicity of the exponential function, means that
$\mathbb{E}_{X\sim\mathbb{Q}}\exp\left(f(X;c)\right)$ $\leq\mathbb{E}_{X\sim\mathbb{Q}}\exp\left(\underbrace{\frac{2\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1}}\right\|_{\mathrm{op}}}{c^{2}}}_{=\frac{1}{c_{1}^{2}}}\frac{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}}{\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}}\right)\stackrel{{\scriptstyle(a)}}{{\leq}}2.$
In $(a)$, we set $c^{2}=2\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1}}\right\|_{\mathrm{op}}c_{1}^{2}$ and used (B.52). ∎
Lemma B.4.
Let $X\sim\mathbb{Q}$, $C_{\mathbb{Q},h_{p}}=\int_{\mathbb{R}^{d}}h_{p}\left(\cdot,{\mathbf{x}}\right)\otimes h_{p}\left(\cdot,{\mathbf{x}}\right)\mathrm{d}\mathbb{Q}(\mathbf{x})$, and $A\in\mathcal{L}(\mathcal{H}_{k})$ a positive operator. Then $\left\|{\left\|{A^{1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}=\operatorname{tr}\left(AC_{\mathbb{Q},h_{p}}\right)$.
Proof.
We have the chain of equalities
@@
$\left\|{\left\|{A^{1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}^{2}\stackrel{{\scriptstyle(a)}}{{=}}\int_{\mathbb{R}^{d}}\left\|{A^{1/2}h_{p}\left(\cdot,{\mathbf{x}}\right)}\right\|_{\mathcal{H}_{k}^{d}}^{2}\mathrm{d}\mathbb{Q}(\mathbf{x})$ (B.57)
=$\stackrel{{\scriptstyle(b)}}{{=}}\int_{\mathbb{R}^{d}}\operatorname{tr}\left[\left(A^{1/2}h_{p}\left(\cdot,{\mathbf{x}}\right)\right)\otimes\left(A^{1/2}h_{p}\left(\cdot,{\mathbf{x}}\right)\right)\right]\mathrm{d}\mathbb{Q}(\mathbf{x})$ (B.58)
=$\stackrel{{\scriptstyle(c)}}{{=}}\operatorname{tr}\left[\int_{\mathbb{R}^{d}}\left(A^{1/2}h_{p}\left(\cdot,{\mathbf{x}}\right)\right)\otimes\left(A^{1/2}h_{p}\left(\cdot,{\mathbf{x}}\right)\right)\mathrm{d}\mathbb{Q}(\mathbf{x})\right]$ (B.59)
=$\stackrel{{\scriptstyle(d)}}{{=}}\operatorname{tr}\left[\int_{\mathbb{R}^{d}}A^{1/2}\left(h_{p}\left(\cdot,{\mathbf{x}}\right)\otimes h_{p}\left(\cdot,{\mathbf{x}}\right)\right)A^{1/2}\mathrm{d}\mathbb{Q}(\mathbf{x})\right]$ (B.60)
=$\stackrel{{\scriptstyle(e)}}{{=}}\operatorname{tr}\left[A^{1/2}\left(\int_{\mathbb{R}^{d}}h_{p}\left(\cdot,{\mathbf{x}}\right)\otimes h_{p}\left(\cdot,{\mathbf{x}}\right)\mathrm{d}\mathbb{Q}(\mathbf{x})\right)A^{1/2}\right]\stackrel{{\scriptstyle(f)}}{{=}}\operatorname{tr}\left(A^{1/2}C_{\mathbb{Q},h_{p}}A^{1/2}\right)$ (B.61)
@@
=$\stackrel{{\scriptstyle(g)}}{{=}}\operatorname{tr}\left(AC_{\mathbb{Q},h_{p}}\right).$ (B.62)
The details are as follows. In (a) we used the definition of $\left\|{\cdot}\right\|_{L_{2}(\mathbb{Q})}$, (b) follows from Lemma D.1 by choosing $f:=A^{1/2}h_{p}\left(\cdot,{\mathbf{x}}\right)$, in (c) the integration was swapped with the $\operatorname{tr}$ [Steinwart and Christmann, 2008, (A.32)]. (d) holds by Lemma B.5 with choosing $a:=b:=h_{p}\left(\cdot,{\mathbf{x}}\right)$, $C:=D:=A^{1/2}$ and using the self-adjointness of $A$ and hence that of $A^{1/2}$. In (e) the integration was flipped with $A^{1/2}$ [Steinwart and Christmann, 2008, (A.32)]. (f) holds by the definition of $C_{\mathbb{Q},h_{p}}$. The cyclic invariance property of the trace implies (g), which concludes the derivation. ∎

### v2

@@ A.3 Proof of Corollary 1
    where (a) uses the exponential decay assumption. (b) uses that $n\geq m$ and that the logarithm is a monotonically increasing function. (c) follows from our choice of $m\gtrsim\sqrt{n}\sqrt{\log\left(1+\frac{c_{1}n}{c\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}\right)\log(12n/\delta)\log(12/\delta)}$, finishing the proof of the corollary.
@@ Appendix B Auxiliary results
This section collects our auxiliary results. Lemma B.1 builds on Rudi et al. [2015, Lemma 6], which assumes bounded feature maps, and on Della Vecchia et al. [2021, Lemma 5], which is stated in the context of leverage scores. Lemma B.2 states that a sub-exponential random variable satisfies Bernstein’s conditions, and Lemma B.3 is about the sub-Gaussianity of norms of Hilbert space-valued random variables. In Lemma B.4, we show how tensor products interplay with linearly transformed vectors. Lemma B.5 is about the maximum of real-valued sub-Gaussian random variables; it is a slightly altered restatement of Canonne [2021].
@@
Let Assumption 1 hold, and assume $0<\lambda\leq\left\|{C_{\mathbb{Q},h_{p}}}\right\|_{\mathrm{op}}$. Then, for any $\delta\in(0,1)$, it holds that
$\left(\mathbb{Q}^{n}\otimes\Lambda^{m}\right)\left(\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\lesssim\lambda\right)\geq 1-\delta$ (B.28)
@@
The next lemma shows that $h_{p}\left(\cdot,{X}\right)$ and the “whitened” random variable $C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)$ enjoy sub-Gaussian properties in terms of their respective $\mathcal{H}_{h_{p}}$ norms.
Lemma B.3 (Sub-Gaussianity of norm of Hilbert space-valued random variables).
Let $\mathcal{H}$ be a separable Hilbert space, $Y\sim\mathbb{Q}\in\mathcal{M}_{1}^{+}(\mathcal{H})$, and $A\in\mathcal{L}(\mathcal{H})$ invertible, self-adjoint, and positive. Assume that $Y$ is sub-Gaussian, in other words $\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}}}\right\|_{\psi_{2}}\lesssim\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}}}\right\|_{L_{2}(\mathbb{Q})}$ holds for all $u\in\mathcal{H}$. Then
$\left\|{\left\|{A^{1/2}Y}\right\|_{\mathcal{H}}}\right\|_{\psi_{2}}^{2}\lesssim\operatorname{tr}\left(A\mathbb{E}_{Y\sim\mathbb{Q}}\left(Y\otimes Y\right)\right).$ (B.45)
Specifically, with Assumption 1, choosing $A:=I$ and $Y:=h_{p}\left(\cdot,{X}\right)$, and $A:=C^{-1}_{\mathbb{Q},h_{p},\lambda}$ ($\lambda>0$) and $Y:=h_{p}\left(\cdot,{X}\right)$, respectively, it holds that
$\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}<\infty,$ and $\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}^{2}\lesssim\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)<\infty,$ (B.46)
that is, both l.h.s. are sub-Gaussian.
@@
Let $\left(e_{i}\right)_{i\in I}$ be a countable ONB of the separable $\mathcal{H}$. We obtain
$\left\|{\left\|{A^{1/2}Y}\right\|_{\mathcal{H}}}\right\|_{\psi_{2}}^{2}\stackrel{{\scriptstyle(a)}}{{=}}\left\|{\left\|{A^{1/2}Y}\right\|_{\mathcal{H}}^{2}}\right\|_{\psi_{1}}\stackrel{{\scriptstyle(b)}}{{=}}\left\|{\sum_{i\in I}\left\langle{A^{1/2}Y,e_{i}}\right\rangle_{\mathcal{H}}^{2}}\right\|_{\psi_{1}}\stackrel{{\scriptstyle(c)}}{{\leq}}\sum_{i\in I}\left\|{\left\langle{A^{1/2}Y,e_{i}}\right\rangle_{\mathcal{H}}^{2}}\right\|_{\psi_{1}}$ (B.47)
@@
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{i\in I}\left\|{\left\langle{A^{1/2}Y,e_{i}}\right\rangle_{\mathcal{H}}}\right\|_{\psi_{2}}^{2}\stackrel{{\scriptstyle(e)}}{{\lesssim}}\sum_{i\in I}\left\|{\left\langle{A^{1/2}Y,e_{i}}\right\rangle_{\mathcal{H}}}\right\|_{L_{2}(\mathbb{Q})}^{2}\stackrel{{\scriptstyle(f)}}{{=}}\sum_{i\in I}\mathbb{E}_{Y\sim\mathbb{Q}}\left\langle{A^{1/2}Y,e_{i}}\right\rangle_{\mathcal{H}}^{2}$ (B.48)
@@
=$\stackrel{{\scriptstyle(g)}}{{=}}\sum_{i\in I}\mathbb{E}_{Y\sim\mathbb{Q}}\left\langle{\left(A^{1/2}Y\right)\otimes\left(A^{1/2}Y\right),e_{i}\otimes e_{i}}\right\rangle_{\mathcal{H}\otimes\mathcal{H}}$ (B.49)
@@
=$\stackrel{{\scriptstyle(h)}}{{=}}\sum_{i\in I}\mathbb{E}_{Y\sim\mathbb{Q}}\left\langle{A^{1/2}\left(Y\otimes Y\right)A^{1/2},e_{i}\otimes e_{i}}\right\rangle_{\mathcal{H}\otimes\mathcal{H}}$ (B.50)
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}\sum_{i\in I}\left\langle{A^{1/2}\mathbb{E}_{Y\sim\mathbb{Q}}\left(Y\otimes Y\right)A^{1/2},e_{i}\otimes e_{i}}\right\rangle_{\mathcal{H}\otimes\mathcal{H}}\stackrel{{\scriptstyle(j)}}{{=}}\sum_{i\in I}\left\langle{A^{1/2}\mathbb{E}_{Y\sim\mathbb{Q}}\left(Y\otimes Y\right)A^{1/2}e_{i},e_{i}}\right\rangle_{\mathcal{H}}$ (B.51)
=$\stackrel{{\scriptstyle(k)}}{{=}}\operatorname{tr}\left(A^{1/2}\mathbb{E}_{Y\sim\mathbb{Q}}\left(Y\otimes Y\right)A^{1/2}\right)\stackrel{{\scriptstyle(l)}}{{=}}\operatorname{tr}\left(A\mathbb{E}_{Y\sim\mathbb{Q}}\left(Y\otimes Y\right)\right).$ (B.52)
The details are as follows. (a) uses Lemma D.2(4), Parseval’s identity yields (b), and the triangle inequality implies (c). (d) holds by Lemma D.2(4). For (e), let $u_{i}=A^{1/2}e_{i}$ and observe that
$\left\|{\left\langle{A^{1/2}Y,e_{i}}\right\rangle_{\mathcal{H}}}\right\|_{\psi_{2}}^{2}\stackrel{{\scriptstyle(m)}}{{=}}\left\|{\left\langle{Y,u_{i}}\right\rangle_{\mathcal{H}}}\right\|_{\psi_{2}}^{2}\stackrel{{\scriptstyle(n)}}{{\lesssim}}\left\|{\left\langle{Y,u_{i}}\right\rangle_{\mathcal{H}}}\right\|_{L_{2}(\mathbb{Q})}^{2}\stackrel{{\scriptstyle(o)}}{{=}}\left\|{\left\langle{A^{1/2}Y,e_{i}}\right\rangle_{\mathcal{H}}}\right\|_{L_{2}(\mathbb{Q})}^{2},$ (B.53)
where (m) uses the self-adjointness of $A^{1/2}$ (implied by the positivity of $A$), (n) follows from the sub-Gaussian assumption on $Y$ holding for arbitrary $u_{i}\in\mathcal{H}$, and (o), again, uses the self-adjointness of $A$. (f) is the definition of the $L_{2}(\mathbb{Q})$-norm, (g) holds by the definition of the tensor product, and Lemma B.4 yields (h). (i) integral and bounded linear operators are swapped by Steinwart and Christmann [2008, (A.32)], (j) is a property of Hilbert-Schmidt operators, and (k) uses the definition of the trace of a linear operator w.r.t. an ONB. The cyclic invariance property of the trace yields (l) and concludes the proof of the first statement.
With $A:=I$ and $Y:=h_{p}\left(\cdot,{X}\right)$, we obtain $\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\lesssim\operatorname{tr}\left(\mathbb{E}\left(h_{p}\left(\cdot,{X}\right)\otimes h_{p}\left(\cdot,{X}\right)\right)\right)=\operatorname{tr}(C_{\mathbb{Q},h_{p}})<\infty$, which is the second statement. The last part follows from considering $A:=C^{-1}_{\mathbb{Q},h_{p},\lambda}$ and $Y:=h_{p}\left(\cdot,{X}\right)$; the invertibility of $C_{\mathbb{Q},h_{p},\lambda}$ guarantees the well-definedness of the $u_{i}$-s ($i\in I$).